A well-known brokerage firm executive claimed that at least 41 % of investors are currently confident of meeting their investment goals. An XYZ Investor Optimism Survey, conducted over a two week period, found that out of 95 randomly selected people, 37 of them said they are confident of meeting their goals. Suppose you are have the following null and alternative hypotheses for a test you are running: H 0 : p = 0.41 H a : p > 0.41 Calculate the test statistic, rounded to 3 decimal places z =
Answer #1

Solution :

Given that,

n = 95 , x = 37

\hat p = x / n = 37 / 95 = 0.3895

P0 = 0.41

1 - P0 = 1 - 0.41 = 0.59

z = \hat p - P0 / [\sqrtP0 * (1 - P0 ) / n]

= 0.3895 - 0.41 / [\sqrt(0.41 * 0.59) / 95]

= -0.407

Test statistic = -0.407